OpenStudy (anonymous):

Write an equation of the line given the following conditions: // to x-3y=-12 & passes through the point (6,5)

OpenStudy (mertsj):

If it is parallel, it has the same slope. The slope of the given line can be found by using the opposite of a divided by b or -1/-3 = 1/3

OpenStudy (mertsj):

Now plug the slope and the given point into point-slope form: y-5=1/3(x-6) Multiply both sides by 3: 3y-15 = x-6 3y=x+9 y=1/3x+3